OpenStudy (anonymous):

Factor completely: x2 − 10x + 25

Parth (parthkohli):

It's a perfect square :D

OpenStudy (anonymous):

^

OpenStudy (anonymous):

Buth there's still an equation right?

Parth (parthkohli):

Yeah, do you know factorization of trinomials?

Parth (parthkohli):

You must express \(b\) as the sum of two numbers \(\alpha \) and \(\beta\) such that \(\alpha \times \beta = a \times c\), first of all.

Parth (parthkohli):

In this case, \(\alpha \) and \(\beta \) are -5 and -5, respectively. \( \color{Black}{\Rightarrow x^2 - 5x - 5x - 10 }\) Did that help?

OpenStudy (anonymous):

Yes thanks .